# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ I.2 Near-Constant Weight Design
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $\geq H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V)$ (403)
$=\sum_{v=0}^{n_{2}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)$ (404)
$\geq\mathbb{P}(\mathscr{A}_{2})H\big{(}\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))\big{)}$ (405)
$\geq(1-o(1))\log\binom{n_{2}}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))},$ (406)
where in (405) we only count the summands in which $V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))$ as specified in the event $\mathscr{A}_{2}$ (402), and (406) holds because given $V=v$ (and $\mathbf{x}_{s_{\mathrm{eq}}}$), $\mathbf{Y}_{2}$ is uniformly distributed over the $\binom{n_{2}}{v}$ possibilities of $n_{2}$-dimensional $\{0,1\}$-valued vector with $v$ $1$’s. Substituting (406) into (394) yields
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\geq(1-o(1))\log\binom{n_{2}}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))}-n_{2}H_{2}(\rho).$ (407)
To evaluate $I_{\ell}^{n}$, we further average over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (389) to obtain that
$I_{\ell}^{n}$ $\geq\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{eq}}}}\big{(}I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\big{)}$ (408)
$\geq(1-o(1))\log\binom{ne^{-\nu(1-\alpha)}(1+o(1))}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))}-\big{(}ne^{-\nu(1-\alpha)}(1+o(1))\big{)}H_{2}(\rho)$ (409)
$=ne^{-\nu(1-\alpha)}(H_{2}(\rho\star e^{-\nu\alpha})-H_{2}(\rho))(1+o(1)),$ (410)
where (409) follows by using the non-negativity of mutual information to sum only over $\mathbf{x}_{s_{\mathrm{eq}}}$ satisfying the event $\mathscr{A}_{1}$ (401) and then using (407) along with $n_{2}=ne^{-\nu(1-\alpha)}(1+o(1))$ (under $\mathscr{A}$), and in (410) we apply $\log{N\choose\beta N}=NH_{2}(\beta)(1+o(1))$ (see Lemma 2). Since this matches the upper bound we derived above, (348) follows.
@@
For the upper bound, we again use (395) to upper bound $H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$, but now trivially use the fact that the probability of a particular test containing an item from $s_{\mathrm{dif}}$ is at most $\frac{\nu\ell}{k}\to 0$ (a union bound of $\frac{\nu n\ell}{k}$ events having probability $\frac{1}{n}$ each). Hence, we have
$\mathbb{P}[Y^{\prime}=1]\leq\frac{\nu\ell}{k}(1-\rho)+\bigg{(}1-\frac{\nu\ell}{k}\bigg{)}\rho=\rho+\frac{\nu\ell}{k}(1-2\rho).$ (411)
@@
$H(Y^{\prime}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq H_{2}\bigg{(}\rho+\frac{\nu\ell}{k}(1-2\rho)\bigg{)}=H_{2}(\rho)+\frac{\nu\ell}{k}(1-2\rho)\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1+o(1)),$ (412)
where we use a Taylor expansion with $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ in the last step. Combining with (394) and (395) gives
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}\frac{\nu\ell}{k}(1-2\rho)\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1+o(1)).$ (413)
Note that $\frac{|s_{\mathrm{eq}}|}{k}=1-\frac{\ell}{k}\to 1$, so (319) in Lemma 7 gives
$\mathbbm{E}(n_{2})=n-\mathbbm{E}(n_{1})=n-\mathbbm{E}(W^{(s_{\mathrm{eq}})})=n-(1-e^{-\nu}+o(1))n=ne^{-\nu}(1+o(1)).$ (414)
Thus, by averaging over $\mathbf{X}_{s_{\mathrm{eq}}}$ in (413), we obtain $I_{\ell}^{n}\leq\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log(\frac{1-\rho}{\rho})(1+o(1))$.
@@
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}_{2},V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (415)
$=H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V)+H(V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (416)
$\geq\sum_{v=0}^{n_{2}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2}),$ (417)
where (415) is because $V$ is deterministic given $\mathbf{Y}_{2}$, (416) holds by the chain rule for entropy [19, Thm. 2.5.1], and (417) holds due to the following argument showing that the entropy of $V$ (when given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$) is at least $\big{(}\frac{1}{2}+o(1)\big{)}\log(n_{2})$:
@@
As before, we seek to identify high-probability events that simplify the further analysis of (417). Fortunately, what we need has already been shown in the proof of Lemma 11, as we now detail.
We first quantify $n_{2}$. By setting $\delta=\Theta\big{(}\sqrt{\frac{\ell}{k}}\big{)}$ in (374) in the proof of Lemma 11, the event
$\mathscr{A}_{3}=\big{\{}n_{2}=ne^{-\nu}(1+o(1))\big{\}}$ (418)
holds with probability at least $1-2\exp(-\frac{C_{*,1}\ell n}{k})$, where $C_{*,1}$ can be chosen arbitrarily large as a function of $(\nu,\rho)$. Similarly, regarding $V$, we set $\delta=\Theta\big{(}\sqrt{\frac{\ell}{k}}\big{)}$ in (376) in the proof of Lemma 11, that the event
$\mathscr{A}_{4}=\big{\{}V=\rho ne^{-\nu}(1+o(1))\big{\}}$ (419)
@@
We argue that it suffices to consider $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$ such that $n_{2}=ne^{-\nu}(1+o(1))$, as given in the high-probability event $\mathscr{A}_{1}$. Recall that our goal is to characterize $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ in (388) and then average it over $\mathbf{X}_{s_{\mathrm{eq}}}$ (as in (389)) to characterize $I_{\ell}^{n}$. Since $\mathbf{Y}$ is binary-valued and has length $n$, we always have $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n\log 2=O(n)$, which means that the total contribution to $I_{\ell}^{n}$ from the low-probability event $\mathscr{A}_{1}^{c}$ is bounded by at most the following:
$O(n)\exp\Big{(}-\frac{C_{*,1}\nu^{-1}\ell n}{k}\Big{)}=\exp\big{(}O(\log k)-C_{*,1}\nu^{-1}\ell\cdot\Omega(\log k)\big{)}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (420)
@@
where the first equality holds since $n=\Theta(k\log k)$, and the second equality follows by taking $C_{*,1}$ to be sufficiently large. Thus, the total contribution under $\mathscr{A}_{1}^{c}$ does not affect the desired bound (347).
@@
For specific $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$ (with $n_{2}=ne^{-\nu}(1+o(1))$ as in the event $\mathscr{A}_{1}$), we define $\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}$ as the corresponding set of $V$ values belonging to the event $\mathscr{A}_{2}$ (all satisfying $V=\rho ne^{-\nu}(1+o(1))$). Then, we obtain from (417) that
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (421)
$\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\log\binom{n_{2}}{v}+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (422)
$\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-\frac{1}{2}\log(n_{2})-\log(2)\Big{]}+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (423)
$=\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}+o\Big{(}\frac{\ell n}{k}\Big{)},$ (424)
@@
where (422) holds because $\mathbf{Y}_{2}|\{V=v\}$ follows a uniform distribution over $\binom{n_{2}}{v}$ possibilities, in (423) we use $\log\binom{N}{\beta N}\geq NH_{2}(\beta)-\frac{1}{2}\log N-\log(2)$ (see Lemma 2), and in (424) we perform some rearrangement to cancel out the terms of $\frac{1}{2}\log(n_{2})$ and then write $o(\log n_{2})=o(\log n)=o(\log k)=o(\frac{\ell n}{k})$, which can be seen analogously to (420). Combining with (394), it follows that
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-n_{2}H_{2}(\rho)\Big{]}$ (425)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\Big{(}\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{)}n_{2}H_{2}(\rho)+o\Big{(}\frac{\ell n}{k}\Big{)}:=\mathcal{T}_{1}-\mathcal{T}_{2}+o\Big{(}\frac{\ell n}{k}\Big{)}.$ (426)
By using $n=\Theta(k\log k)$ and choosing $C_{*,2}$ large enough, analogously to (420), we obtain
$|\mathcal{T}_{2}|\leq\mathbb{P}(\mathscr{A}_{2}^{c})n_{2}H_{2}(\rho)$ $\leq 2\exp\Big{(}-\frac{C_{*,2}\ell n}{k}\Big{)}nH_{2}(\rho)$ (427)
$=O\Big{(}\exp\big{(}O(\log k)-C_{*,2}\ell\cdot\Omega(\log k)\big{)}\Big{)}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (428)
@@
which only amounts to a lower-order term in the desired result (347).
@@
Thus, by further averaging over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (389), we only need to prove $\mathbbm{E}(\mathcal{T}_{1})$ is lower bounded by the right-hand side of (347). Regarding $\mathcal{T}_{1}$, since we are considering $n_{2}=ne^{-\nu}(1+o(1))$ and $v=\rho ne^{-\nu}(1+o(1))$ (see the high-probability events in (401)–(402)), a Taylor expansion with derivative $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ gives
$H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-H_{2}(\rho)=\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\Big{(}\frac{v}{n_{2}}-\rho\Big{)}(1+o(1)).$ (429)
@@
We now substitute this into $\mathcal{T}_{1}$, then by decomposing $\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}=\sum_{v=0}^{n_{2}}-\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}$ we can write $\mathcal{T}_{1}$ as

### 23 Oct 2024 v2

@@ I.2 Near-Constant Weight Design
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $\geq H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V)$ (J.52)
$=\sum_{v=0}^{n_{2}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)$ (J.53)
$\geq\mathbb{P}(\mathscr{A}_{2})H\big{(}\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))\big{)}$ (J.54)
$\geq(1-o(1))\log\binom{n_{2}}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))},$ (J.55)
where in (J.54) we only count the summands in which $V=ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))$ as specified in the event $\mathscr{A}_{2}$ (J.51), and (J.55) holds because given $V=v$ (and $\mathbf{x}_{s_{\mathrm{eq}}}$), $\mathbf{Y}_{2}$ is uniformly distributed over the $\binom{n_{2}}{v}$ possibilities of $n_{2}$-dimensional $\{0,1\}$-valued vector with $v$ $1$’s. Substituting (J.55) into (J.43) yields
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\geq(1-o(1))\log\binom{n_{2}}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))}-n_{2}H_{2}(\rho).$ (J.56)
To evaluate $I_{\ell}^{n}$, we further average over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (J.38) to obtain that
$I_{\ell}^{n}$ $\geq\mathbbm{E}_{\mathbf{X}_{s_{\mathrm{eq}}}}\big{(}I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\big{)}$ (J.57)
$\geq(1-o(1))\log\binom{ne^{-\nu(1-\alpha)}(1+o(1))}{ne^{-\nu(1-\alpha)}(\rho\star e^{-\nu\alpha})(1+o(1))}-\big{(}ne^{-\nu(1-\alpha)}(1+o(1))\big{)}H_{2}(\rho)$ (J.58)
$=ne^{-\nu(1-\alpha)}(H_{2}(\rho\star e^{-\nu\alpha})-H_{2}(\rho))(1+o(1)),$ (J.59)
where (J.58) follows by using the non-negativity of mutual information to sum only over $\mathbf{x}_{s_{\mathrm{eq}}}$ satisfying the event $\mathscr{A}_{1}$ (J.50) and then using (J.56) along with $n_{2}=ne^{-\nu(1-\alpha)}(1+o(1))$ (under $\mathscr{A}$), and in (J.59) we apply $\log{N\choose\beta N}=NH_{2}(\beta)(1+o(1))$ (see Lemma A.2). Since this matches the upper bound we derived above, (H.9) follows.
@@
In order to upper bound $H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$, we again use (J.44), but we now trivially use the fact that the probability of a particular test containing an item from $s_{\mathrm{dif}}$ is at most $\frac{\nu\ell}{k}\to 0$ (a union bound of $\frac{\nu n\ell}{k}$ events having probability $\frac{1}{n}$ each). Hence, we have
$\mathbb{P}[Y^{\prime}=1]\leq\frac{\nu\ell}{k}(1-\rho)+\bigg{(}1-\frac{\nu\ell}{k}\bigg{)}\rho=\rho+\frac{\nu\ell}{k}(1-2\rho).$ (J.60)
@@
$H(Y^{\prime}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq H_{2}\bigg{(}\rho+\frac{\nu\ell}{k}(1-2\rho)\bigg{)}=H_{2}(\rho)+\frac{\nu\ell}{k}(1-2\rho)\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1+o(1)),$ (J.61)
where we use a Taylor expansion with $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ in the last step. Combining with (J.43) and (J.44) gives
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n_{2}\frac{\nu\ell}{k}(1-2\rho)\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1+o(1)).$ (J.62)
Note that $\frac{|s_{\mathrm{eq}}|}{k}=1-\frac{\ell}{k}\to 1$, so (G.12) in Lemma G.3 gives
$\mathbbm{E}(n_{2})=n-\mathbbm{E}(n_{1})=n-\mathbbm{E}(W^{(s_{\mathrm{eq}})})=n-(1-e^{-\nu}+o(1))n=ne^{-\nu}(1+o(1)).$ (J.63)
Thus, by averaging over $\mathbf{X}_{s_{\mathrm{eq}}}$ in (J.62), we obtain $I_{\ell}^{n}\leq\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log(\frac{1-\rho}{\rho})(1+o(1))$.
@@
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $=H(\mathbf{Y}_{2},V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (J.64)
$=H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V)+H(V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ (J.65)
$\geq\sum_{v=0}^{n_{2}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2}),$ (J.66)
where (J.64) is because $V$ is deterministic given $\mathbf{Y}_{2}$, (J.65) holds by the chain rule for entropy [19, Thm. 2.5.1], and (J.66) holds due to the following argument showing that the entropy of $V$ (when given $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$) is at least $\big{(}\frac{1}{2}+o(1)\big{)}\log(n_{2})$:
@@
As before, we seek to identify high-probability events that simplify the further analysis of (J.66). Fortunately, what we need has already been shown in the proof of Lemma H.3, as we now detail.
We first quantify $n_{2}$. By setting $\delta=\Theta\big{(}\sqrt{\frac{\ell}{k}}\big{)}$ in (J.23) in the proof of Lemma H.3, the event
$\mathscr{A}_{3}=\big{\{}n_{2}=ne^{-\nu}(1+o(1))\big{\}}$ (J.67)
holds with probability at least $1-2\exp(-\frac{C_{*,1}\ell n}{k})$, where $C_{*,1}$ can be chosen arbitrarily large as a function of $(\nu,\rho)$. Similarly, regarding $V$, we set $\delta=\Theta\big{(}\sqrt{\frac{\ell}{k}}\big{)}$ in (J.25) in the proof of Lemma H.3, that the event
$\mathscr{A}_{4}=\big{\{}V=\rho ne^{-\nu}(1+o(1))\big{\}}$ (J.68)
@@
We argue that it suffices to consider $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$ such that $n_{2}=ne^{-\nu}(1+o(1))$, as given in the high-probability event $\mathscr{A}_{1}$. Recall that our goal is to characterize $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ in (J.37) and then average it over $\mathbf{X}_{s_{\mathrm{eq}}}$ (as in (J.38)) to characterize $I_{\ell}^{n}$. Since $\mathbf{Y}$ is binary-valued and has length $n$, we always have $I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\leq n\log 2=O(n)$, which means that the total contribution to $I_{\ell}^{n}$ from the low-probability event $\mathscr{A}_{1}^{c}$ is bounded by at most the following:
$O(n)\exp\Big{(}-\frac{C_{*,1}\nu^{-1}\ell n}{k}\Big{)}=\exp\big{(}O(\log k)-C_{*,1}\nu^{-1}\ell\cdot\Omega(\log k)\big{)}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (J.69)
@@
where the first equality holds since $n=\Theta(k\log k)$ (see Section A.2), and the second equality follows by taking $C_{*,1}$ to be sufficiently large. Thus, the total contribution under $\mathscr{A}_{1}^{c}$ does not affect the desired bound (H.8).
@@
For specific $\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}$ (with $n_{2}=ne^{-\nu}(1+o(1))$ as in the event $\mathscr{A}_{1}$), we define $\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}$ as the corresponding set of $V$ values belonging to the event $\mathscr{A}_{2}$ (all satisfying $V=\rho ne^{-\nu}(1+o(1))$). Then, we obtain from (J.66) that
$H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})$ $\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)H(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},V=v)+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (J.70)
$\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\log\binom{n_{2}}{v}+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (J.71)
$\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-\frac{1}{2}\log(n_{2})-\log(2)\Big{]}+\Big{(}\frac{1}{2}+o(1)\Big{)}\log(n_{2})$ (J.72)
$=\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}+o\Big{(}\frac{\ell n}{k}\Big{)},$ (J.73)
@@
where (J.71) holds because $\mathbf{Y}_{2}|\{V=v\}$ follows a uniform distribution over $\binom{n_{2}}{v}$ possibilities, in (J.72) we use $\log\binom{N}{\beta N}\geq NH_{2}(\beta)-\frac{1}{2}\log N-\log(2)$ (see Lemma A.2), and in (J.73) we perform some rearrangement to cancel out the terms of $\frac{1}{2}\log(n_{2})$ and then write $o(\log n_{2})=o(\log n)=o(\log k)=o(\frac{\ell n}{k})$, which can be seen analogously to (J.69). Combining with (J.43), it follows that
$I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}})\geq\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}n_{2}H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-n_{2}H_{2}(\rho)\Big{]}$ (J.74)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\Big{(}\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{)}n_{2}H_{2}(\rho)+o\Big{(}\frac{\ell n}{k}\Big{)}:=\mathcal{T}_{1}-\mathcal{T}_{2}+o\Big{(}\frac{\ell n}{k}\Big{)}.$ (J.75)
By using $n=\Theta(k\log k)$ and choosing $C_{*,2}$ large enough, analogously to (J.69), we obtain
$|\mathcal{T}_{2}|\leq\mathbb{P}(\mathscr{A}_{2}^{c})n_{2}H_{2}(\rho)$ $\leq 2\exp\Big{(}-\frac{C_{*,2}\ell n}{k}\Big{)}nH_{2}(\rho)$ (J.76)
$=O\Big{(}\exp\big{(}O(\log k)-C_{*,2}\ell\cdot\Omega(\log k)\big{)}\Big{)}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (J.77)
@@
which only amounts to a lower-order term in the desired result (H.8).
@@
Thus, by further averaging over $\mathbf{X}_{s_{\mathrm{eq}}}$ as in (J.38), we only need to prove $\mathbbm{E}(\mathcal{T}_{1})$ is lower bounded by the right-hand side of (H.8). Since we are considering $n_{2}=ne^{-\nu}(1+o(1))$ and $v=\rho ne^{-\nu}(1+o(1))$ (see the high-probability events in (J.50)–(J.51)), a Taylor expansion with derivative $H^{\prime}_{2}(\rho)=\log\frac{1-\rho}{\rho}$ gives
$H_{2}\Big{(}\frac{v}{n_{2}}\Big{)}-H_{2}(\rho)=\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\Big{(}\frac{v}{n_{2}}-\rho\Big{)}(1+o(1)).$ (J.78)
@@
We substitute this into $\mathcal{T}_{1}$, then decompose $\sum_{v\in\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}=\sum_{v=0}^{n_{2}}-\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}$ to write $\mathcal{T}_{1}$ as